If you have a grid of 24 squares, use the grid alongside to illustrate 3/8 and 2/3. Show the answer to 2/3-3/8 Clearly by making use of different types of shading


1
Expert's answer
2021-05-21T03:46:03-0400

Solution:

"B=\frac38=\frac38\times \frac33=\frac9{24}
\\ A=\frac23=\frac23\times \frac88=\frac{16}{24}"



Now, on subtracting B from A, we get C, which denotes "\frac7{24}"